LHC Report: finalizing the shutdown activities

The maintenance work and other activities scheduled for the LHC technical stop have now been completed and the electrical, quality assurance and powering tests are in full swing.


These hardware tests, or hardware re-commissioning as it is known in the CERN Control Centre, are complete for Sectors 5-6 and 6-7. The re-commissioning process is almost complete in Sectors 7-8 and 8-1, but a problem with the emergency stop safety system last week, and the failure of a turbine in the cryogenic plant at Point 8, mean that the final part of the re-commissioning for these two sectors has been delayed and will be completed this week. Preparations for the re-commissioning in the other 4 sectors are going well, and everything is on schedule for the LHC to restart with beam as planned on 18 February.

At the SPS, all the technical stop work and magnet changes have been completed and the machine has been handed over to the Operations Group for the usual set of hardware tests and preparations for beam operation.

Finally all the technical stop activities at the PS are complete and the machine has been given to TE/EPC for the initial commissioning of the new POPS power supply, which will replace the venerable old MPS. Last Thursday, POPS successfully powered the PS magnets for the first time up to the nominal current.

The POPS performance tests are taking place this week with the involvement of the CCC, initially without beam. Then, if the results are conclusive, the tests will begin with beam on Friday.

by CERN Bulletin